# Likelihood-based Sensor Calibration for Expert-Supported Distributed Learning Algorithms in IoT Systems

### v1

@@ Abstract
An important task in the field of sensor technology is the efficient implementation of adaptation procedures of measurements from one sensor to another sensor of identical design. One idea is to use the estimation of an affine transformation between different systems, which can be improved by the knowledge of experts. This paper presents an improved solution from Glacier Research that was published back in 1973. It is shown that this solution can be adapted for software calibration of sensors, implementation of expert-based adaptation, and federated learning methods. We evaluate our research with simulations and also with real measured data of a multi-sensor board with 8 identical sensors. The results show an improvement for both the simulation and the experiments with real data.
@@ I Introduction
By combining these two methods, we lay the foundation for improving transformation estimation through expert knowledge. This can be used in federated learning systems by checking the computed transformation between similar systems by an expert, and iteratively triggering a recomputation or further adaptation of the transformation in case of incorrect predictions. The given example of 8 sensors measuring the same fluid, will provide similar data as if 8 sensors were measuring 8 similar fluids at different locations, which would correspond to a distributed learning use case.
@@
(a) Raw measurement data of an experiment with an artificial nose.
(b) Normalized measurement data of an experiment with an artificial nose.
@@
Inspired by a statistical problem in glaciology, Gleser et al. [1] obtained maximum likelihood estimators (MLE) of the unknown features and an AT. As a continuation of this research, the authors in [1] showed that a MLE for the unknown parameter exists, and gained MLE for the model parameters in closed form.
@@ I-A Related Work
More general approach for the MLE procedure for linear model with errors in variables is proposed by authors in [6, 8]. In [9], the relationship between maximum likelihood and generalize least squares approaches are examined and some of the results when estimating in error in variables models are extended. The linear transformations trained in maximum likelihood sense on adaptation data for hidden Markov model based speech recognition is researched in [10, 11]. The authors in [12] discuss and compare the optimal linear estimation and modified MLE, where they show the better results for modified MLE from the point of bias and means square error in comparison with optimal linear model. The maximum likelihood method is used in regression analysis of linear transformation model and efficient and consistent estimator is developed for different data case of cohorts [13, 14]. MLE optical flow model in monocular optical flow field is developed and presented by authors in [15]. Domain adaptation using maximum likelihood linear transformation for speaker verification is presented by authors in [16, 17]. The authors in [19] investigate the transformed linear regression with non-normal error distributions where they develop a MLE approach and provide the almost optimal conditions on the error distribution. The authors in [20], by using Student’s t-distribution noise model, introduce a distributed maximum-likelihood based state estimation approach in domain of power systems. The authors in [18] propose a Federated Learning framework robust to affine distribution shifts (FLRA) for non-i.i.d. personalization to implement efficient federated learning to generate an averaged global model using ATs. In [22], AT parameters are used to analyze the domain heterogeneity for human activity recognition.
I-B Contribution
In this paper, we show that the results of an old but fundamental work [1] in glaciology can be used for the estimation of an AT in expert-supported learning. In [1], the authors derive a solution for the estimation of a linear transformation. Their solution allows to estimate the data in system 1, variance and the transformation. By using the well-known augmented matrix, this result can be easily improved by denoising using the eigenvectors and extended to an AT. Furthermore, the authors already showed that a simple solution directly follows from the multivariate regression analysis. We are interested in the estimation of the AT and we additionally derive the gradient of the MLE to get a stationary point of the parameters of the transformation, which leads to a similar simplificated least square solution as in [1]. We prove that both solutions are connected by an eigenvalue decomposition. By simulation, we show that a combination of both solutions provides better estimates for computing the parameters of the AT compared to the original approach from Gleser et al. [1]. We also show an experiment with a multi-sensor Bosch BME688 development kit board (figure 3) to evaluate the estimation error of the different methods and provide the example dataset and evaluation code. Finally, we present a novel concept for distributed learning using transformation with the support of expert knowledge.
I-C Organization of the Paper
In section II, we derive the system model and formulate the estimation problem of the AT. In section III, we present the algorithms, which we in part have selected from the literature [1]. Here in section III-A, we summarize the solution of Gleser et al.. We show in section III-B that the stationary points of the parameters of the transformation lead to similar results as the least squares solution, which can be combined with Gleser et al. to provide better estimation of the AT. Section IV summarizes the simulation results based on Monte Carlo simulations and shows the results in a real multi-sensor setup.
@@ II System Model
In this section, we define the mathematical basis in order to develop the corresponding algorithms in section III. Figure 4 shows an example of the transformation for the 2-dimensional case. This figure presents a simulation of the data vectors $\mathbf{v}\in\mathbb{R}^{2}$ in an Euclidean space. Here, the black data points are the measured points $\mathbf{X}$ from system $1$ and the red data points $\boldsymbol{\Theta}$ are the estimation of these data points. The green data points $\mathbf{Y}$ represent the (noisy) transformation of the black data points $\mathbf{X}$ into system $2$ and the blue data points $\hat{\mathbf{Y}}$ are the origin estimation of this transformed data.
As in [1], we consider two different systems in which the same features are measured. A single measurement data vector of system $1$ is denoted by $\mathbf{x}_{i}$ and the single data measurement vector of system $2$ by $\mathbf{y}_{i}$ respectively, where $i\in\{1,...,n\}$. Let $\mathbf{x}_{i}\in\mathbb{R}^{q}$ and $\mathbf{y}_{i}\in\mathbb{R}^{q}$, then the data vectors from system $1$ are mapped to the data vectors in system $2$ by the transformation $\mathbf{y}_{i}=\mathrm{T}(\mathbf{x}_{i})$. We can summarize in the following definition:
@@
We assume $\mathbf{m}_{i}$ and $\mathbf{n}_{i}$ are statistically independent. The covariance matrix of $\mathbf{n}_{i}$ and $\mathbf{m}_{i}$ is $\mathrm{Cov}[\mathbf{m}_{i}]=\mathrm{Cov}[\mathbf{n}_{i}]=\sigma^{2}\mathbf{I}$.
We can summarize all measurements to the measurement matrices $\mathbf{X}_{M}=[\mathbf{x}_{1},...,\mathbf{x}_{n}]$, and $\mathbf{Y}_{M}=[\mathbf{y}_{1},...,\mathbf{y}_{n}]$, and the estimated origin matrix $\boldsymbol{\Theta}_{E}=[\boldsymbol{\theta}_{1},...,\boldsymbol{\theta}_{n}]$.
@@ III Algorithms
The solution of Gleser et al. [1] was only derived for a linear transformation without a translation vector $\mathbf{b}$. To introduce the estimation with the translation vector $\mathbf{b}$, we extend our input using the well known definition of the augmented matrix in (4)
@@
and the new augmented vectors $\hat{\mathbf{y}}_{i}=[\mathbf{y}_{i}^{T},1]^{T}$, $\hat{\mathbf{x}}_{i}=[\mathbf{x}_{i}^{T},1]^{T}$, $\hat{\boldsymbol{\theta}}_{i}=[\boldsymbol{\theta}_{i}^{T},1]^{T}$, $\hat{\mathbf{m}_{i}}=[\mathbf{m}_{i}^{T},0]^{T}$, and $\hat{\mathbf{n}_{i}}=[\mathbf{n}_{i}^{T},0]^{T}$ the estimation problem of (3) and (1) can be represented by equations (5) and (6):
@@
III-A Solution of Gleser et al.
@@
[1] Let $\boldsymbol{\lambda}$ be the diagonal matrix of the $p$ largest eigenvalues and $\mathbf{U}$ be the matrix of all corresponding eigenvectors of $\mathbf{X}^{T}\mathbf{X}+\mathbf{Y}^{T}\mathbf{Y}$, then the estimates of $\boldsymbol{\Theta}$ and $\mathbf{B}$ are given by (9) and (10):
@@ III-B Multivariate Regression Approach
In our approach, we are not focussing on $\boldsymbol{\Theta}$. We are interested in an estimation of the AT from $\mathbf{X}$ to $\mathbf{Y}$. We can rewrite (6) to  (11):
@@
Then the covariance matrix of $\mathbf{r}_{i}=\hat{\mathbf{n}}_{i}-\mathbf{B}\hat{\mathbf{m}}_{i}$ is given by (12):
$\mathrm{Cov}[\mathbf{r}_{i}]$ $=\mathrm{Cov}[\hat{\mathbf{n}}_{i}-\mathbf{B}\hat{\mathbf{m}}_{i}]$
$=\mathrm{Cov}[\hat{\mathbf{n}}_{i}]+\mathrm{Cov}[\mathbf{B}\hat{\mathbf{m}}_{i}]$
$=\sigma^{2}(\mathbf{I}+\mathbf{B}\mathbf{B}^{T}).$ (12)
In real sensor measurements, we measure $\mathbf{X}$ only, therefore a simple estimation based on $\mathbf{\Theta}=\mathbf{X}$ can be directly used due to the expectation of the noise $m_{i}$ and $n_{i}$ is zero. However, then we will have no estimation of the origin $\mathbf{\Theta}$ of the measured data $\mathbf{X}$. Usually the sensor measurement is also noisy. However, in cases of very low noise power, this approach has advantages, e.g., lower computational and thus energy requirements. The result of this approach is the simple least squares estimate [1]. In the evaluation, we also considered this idea in the (18) and Alg. 2 and 3 for comparison.
As observed in [1], if $\boldsymbol{\Theta}\boldsymbol{\Theta}^{T}$ is non-singluar, the minimum over $\mathbf{B}$ of $\mathrm{tr}[(\mathbf{Y}-\mathbf{B}\boldsymbol{\Theta})(\mathbf{Y}-\mathbf{B}\boldsymbol{\Theta})^{T}]$ can be obtained based on the results of multivariate regression $\mathbf{B}=\mathbf{Y}\boldsymbol{\Theta}^{T}[\boldsymbol{\Theta}\boldsymbol{\Theta}^{T}]^{-1}$. This solution corresponds to the search for $\mathbf{B}$ with $\mathbf{B}\boldsymbol{\Theta}=\mathbf{Y}$. The resulting matrix $\mathbf{B}$ is obtained by multiplying $\mathbf{Y}$ by the pseudoinverse of $\boldsymbol{\Theta}$ (i. e., the Moore-Penrose inverse), i.e. $\mathbf{B}=\mathbf{Y}\boldsymbol{\Theta}^{T}[\boldsymbol{\Theta}\boldsymbol{\Theta}^{T}]^{-1}$. This solution is the least squares solution.
Interestingly, the stationary points of (8) are $\boldsymbol{\Theta}=\mathbf{X}$ and $\mathbf{Y}=\mathbf{B}\boldsymbol{\Theta}$.
Figure 5: Simulation result, mean errors $\bar{e}_{x}$ and $\bar{e}_{y}$ for the extended implementation of Gleser et al. compared to the extended Alg. 1 as a function of the standard deviation $\sigma$.
@@
Algorithm 1 Augmented implementation of Gleser et al. (denoised)
@@
[…]
2:  $[\mathbf{U},\boldsymbol{\Lambda}]=\mathrm{eigs}(\mathbf{X}^{T}\mathbf{X}+\mathbf{Y}^{T}\mathbf{Y},p)$
3:  $\boldsymbol{\Theta}^{T}=\mathbf{U}\mathbf{U}^{T}\mathbf{X}$
4:  $\boldsymbol{\Theta}(p,:)=\mathbf{1}^{T}$
5:  $\mathbf{B}=\mathbf{Y}\boldsymbol{\Theta}^{T}(\boldsymbol{\Theta}\boldsymbol{\Theta}^{T})^{-1}$
6:  $\tilde{\mathbf{A}}=\mathbf{B}(1:p-1,1:p-1)$
7:  $\tilde{\mathbf{b}}=\mathbf{b}(1:p-1,p)$
8:  $\boldsymbol{\Theta}_{E}=\boldsymbol{\Theta}(1:p-1,:)$;
We can now summarize the first two algorithms. Algorithm 1 is the augmented implementation of Gleser et al.. In line $2$ we calculate the $p$ largest eigenvalues $\boldsymbol{\Lambda}$ according to Lemma 1, and use the associated eigenvectors to estimate $\boldsymbol{\Theta}$, the origins of $\mathbf{X}$, in line $3$. The resulting $\boldsymbol{\Theta}$ augmentation line contains many values close to $1$, so we set those to the value $1$ in line $4$ to suppress the small eigenvalues impact on $\boldsymbol{\Theta}$ with low noise power [24]. Lines $5$ to $7$ calculate and extract the estimated AT parameters and line $8$ removes the augmentation to format the origin estimation.
Algorithm 2 corresponds to the derivation of the simple least squares according to Proposition 1. Since our focus lies not in the estimation of the origins $\boldsymbol{\Theta}$, we can directly calculate and extract the estimatation parameters in lines $2$ to $5$. Although we cannot achieve improved complexity, it should be possible to achieve lower energy consumption through an improved number of operations, since Alg. 2 skips step 2 and step 3 of Alg. 1. Furthermore, the accuracy is very close to the results of Alg. 3 (cf. table I and table II), which, however, requires a similar number of operations as Alg. 1.
@@
[…]
2:  $\mathbf{B}=\mathbf{Y}\mathbf{X}^{T}(\mathbf{X}\mathbf{X}^{T})^{-1}$
3:  $\tilde{\mathbf{A}}=\mathbf{B}(1:p-1,1:p-1)$
4:  $\tilde{\mathbf{b}}=\mathbf{b}(1:p-1,p)$
5:  $\boldsymbol{\Theta}_{E}=\mathbf{X}(1:p-1,:)$;
TABLE I: Averaged simulation errors $\bar{e}_{x}$ and $\bar{e}_{y}$ for implementation of Gleser et al., Alg. 1, Alg. 2, and Alg. 3 with increasing standard deviation $\sigma$.
Algorithm: Standard deviation $\sigma$: 1 3 5 7 9 11 13 15
Gleser et al. $\bar{e}_{x}$ $\mathbf{1.0975}$ $\mathbf{3.2894}$ $\mathbf{5.4887}$ $\mathbf{7.6816}$ $\mathbf{9.8777}$ $\mathbf{12.0760}$ $\mathbf{14.2788}$ $\mathbf{16.4682}$
Gleser et al. $\bar{e}_{y}$ $0.6479$ $2.0236$ $3.5964$ $5.4476$ $7.5847$ $10.0661$ $12.8498$ $15.9657$
The simple least squares solution of Alg. 2 and the solution of Gleser et al. in Alg. 1 can be combined to the improved Alg. 3 by using the following observations.

### v4

# Likelihood-based Sensor Calibration using Affine Transformation

@@ Abstract
An important task in the field of sensor technology is the efficient implementation of adaptation procedures of measurements from one sensor to another sensor of identical design. One idea is to use the estimation of an affine transformation between different systems, which can be improved by the knowledge of experts. This paper presents an improved solution from Glacier Research that was published back in 1973. The results demonstrate the adaptability of this solution for various applications, including software calibration of sensors, implementation of expert-based adaptation, and paving the way for future advancements such as distributed learning methods. One idea here is to use the knowledge of experts for estimating an affine transformation between different systems. We evaluate our research with simulations and also with real measured data of a multi-sensor board with 8 identical sensors. Both data set and evaluation script are provided for download. The results show an improvement for both the simulation and the experiments with real data.
@@ I Introduction
By combining these two methods, we lay the foundation for improving transformation estimation through expert knowledge. This can be used in distributed learning systems by checking the computed transformation between similar systems by an expert, and iteratively triggering a recomputation or further adaptation of the transformation in case of incorrect predictions. The given example of 8 sensors measuring the same fluid, will provide similar data as if 8 sensors were measuring 8 similar fluids at different locations, which would correspond to a distributed learning use case.
@@
(a) Raw measurement data of an experiment with an electronic nose.
(b) Normalized measurement data of an experiment with an electronic nose.
@@
Inspired by a statistical problem in glaciology, Gleser and Watson [1] obtained maximum likelihood estimators (MLE) of the unknown features and an AT. As a continuation of this research, the authors in [1] showed that a MLE for the unknown parameter exists, and gained MLE for the model parameters in closed form.
This paper is organized as follows: in section II, we derive the system model and formulate the estimation problem of the AT. In section III, we present the algorithms, which we in part have selected from the literature [1]. Section III-A summarizes the solution of Gleser and Watson. We show in section III-C that the stationary points of the parameters of the transformation lead to similar results as the least squares solution, which can be combined with Gleser and Watson to provide better estimation of the AT. Section IV summarizes the simulation results based on Monte Carlo simulations and shows the results in a real multi-sensor setup.
Our main contribution is outlined as follows:
  1. 1.
    We show that the results of an old but fundamental work [1] in glaciology can be used for the estimation of an AT in expert-supported learning. In [1], the authors derive a solution for the estimation of a linear transformation. Their solution allows to estimate the data in system 1, variance and the transformation. By using the well-known augmented matrix [2], this result can be easily improved by denoising using the eigenvectors and extended to an AT.
  2. 2.
    Furthermore, the authors already showed that a simple solution directly follows from the multivariate regression analysis. We are interested in the estimation of the AT and we additionally derive the gradient of the MLE to get a stationary point of the parameters of the transformation, which leads to a similar simplificated least square solution as in [1].
  3. 3.
    We prove that both solutions are connected by an eigenvalue decomposition. By simulation, we show that a combination of both solutions provides better estimates for computing the parameters of the AT compared to the original approach from Gleser and Watson [1].
  4. 4.
    We also show an experiment with a multi-sensor Bosch BME688 development kit board (Fig. 3) to evaluate the estimation error of the different methods and provide the example dataset and evaluation code.
  5. 5.
    Finally, we briefly present an idea for a novel concept for distributed learning using transformation with the support of expert knowledge.
@@ I-A Related Work
More general approach for the MLE procedure for linear model with errors in variables is proposed by authors in [7, 9]. In [10], the relationship between maximum likelihood and generalize least squares approaches are examined and some of the results when estimating in error in variables models are extended. The linear transformations trained in maximum likelihood sense on adaptation data for hidden Markov model based speech recognition is researched in [11, 12]. The authors in [13] discuss and compare the optimal linear estimation and modified MLE, where they show the better results for modified MLE from the point of bias and means square error in comparison with optimal linear model. The maximum likelihood method is used in regression analysis of linear transformation model and efficient and consistent estimator is developed for different data case of cohorts [14, 15]. MLE optical flow model in monocular optical flow field is developed and presented by authors in [16].
Domain adaptation using maximum likelihood linear transformation for speaker verification is presented by authors in [17, 18]. The authors in [20] investigate the transformed linear regression with non-normal error distributions where they develop a MLE approach and provide the almost optimal conditions on the error distribution. The authors in [21], by using Student’s t-distribution noise model, introduce a distributed maximum-likelihood based state estimation approach in domain of power systems. The authors in [19] propose a Distributed Learning framework robust to affine distribution shifts (FLRA) for non-i.i.d. personalization to implement efficient fistributed learning to generate an averaged global model using ATs. In [23], AT parameters are used to analyze the domain heterogeneity for human activity recognition. The research by [28] introduces a cost-effective calibration method for Low Cost Air Quality Monitoring Systems, which ensures accurate measurements for widespread deployment. This approach reduces calibration costs and enhances the accessibility of precise readings, making it suitable for various IoT Air Quality monitoring devices. The authors in [29] propose TDACNN, a target-domain-free domain adaptation convolutional neural network designed to mitigate the effects of ill-posed gas sensor drift in E-nose applications. The introduction of an additive angular margin softmax loss during training and the use of the MMD-based ensemble method contribute to improved feature generalization and interclass distance enlargement without the need for participation of target domain data. The authors in [30] built five twin sensing units with eight MOX gas sensors each, exploring sensor tolerance and the impact of sensor mismatch, and they demonstrated the effectiveness of the Direct Standardization (DS) transformation for calibration transfer, addressing drift and variability issues. Tao et al. [31] propose a domain correction based on the kernel transformation (DCKT) for drift compensation. They show that transformation can lead to better classification results in drift-prone sensor setups by modifying source and target domain. Authors in [32, 33] addresses flaws in Maximum Likehood Calibration Array (MLCA), proposing a modified cost function and emphasizing improved convergence rates, while our proposed approach focuses on expert-assisted learning between systems, combining direct estimation with a solution from Gleser and Watson.
@@ II System Model
In this section, we define the mathematical basis in order to develop the corresponding algorithms in Section III. Fig. 4 shows an example of the transformation for the 2-dimensional case. This figure presents a simulation of the data vectors $\mathbf{v}\in\mathbb{R}^{2}$ in an Euclidean space. The black data points are the measured points $\mathbf{X}$ from system $1$ and the red data points $\boldsymbol{\Theta}$ are the estimation of these data points. The green data points $\mathbf{Y}$ represent the (noisy) transformation of the black data points $\mathbf{X}$ into system $2$ and the blue data points $\hat{\mathbf{Y}}$ are the origin estimation of this transformed data. As mentioned in [1], we consider two different systems in which the same features are measured. A single measurement data vector of system $1$ is denoted by $\mathbf{x}_{i}$ and the single data measurement vector of system $2$ by $\mathbf{y}_{i}$ respectively, where $i\in\{1,...,n\}$. Let $\mathbf{x}_{i}\in\mathbb{R}^{q}$ and $\mathbf{y}_{i}\in\mathbb{R}^{q}$, then the data vectors from system $1$ are mapped to the data vectors in system $2$ by the transformation $\mathbf{y}_{i}=\mathrm{T}(\mathbf{x}_{i})$.
We can summarize in the following definition:
@@
We assume $\mathbf{m}_{i}$ and $\mathbf{n}_{i}$ are statistically independent. The covariance matrix of $\mathbf{n}_{i}$ and $\mathbf{m}_{i}$ is $\mathrm{Cov}[\mathbf{m}_{i}]=\mathrm{Cov}[\mathbf{n}_{i}]=\sigma^{2}\mathbf{I}$. We can summarize all measurements to the measurement matrices $\mathbf{X}_{M}=[\mathbf{x}_{1},...,\mathbf{x}_{n}]$, and $\mathbf{Y}_{M}=[\mathbf{y}_{1},...,\mathbf{y}_{n}]$, and the estimated origin matrix $\boldsymbol{\Theta}_{E}=[\boldsymbol{\theta}_{1},...,\boldsymbol{\theta}_{n}]$.
@@ III Algorithms
The solution of Gleser and Watson [1] was only derived for a linear transformation without a translation vector $\mathbf{b}$. To introduce the estimation with the translation vector $\mathbf{b}$, we extend our input using the well known definition of the augmented matrix in (4)
@@
and the new augmented vectors $\hat{\mathbf{y}}_{i}=[\mathbf{y}_{i}^{T},\beta_{i}]^{T}$, $\hat{\mathbf{x}}_{i}=[\mathbf{x}_{i}^{T},\alpha_{i}]^{T}$, $\hat{\boldsymbol{\theta}}_{i}=[\boldsymbol{\theta}_{i}^{T},\gamma_{i}]^{T}$, $\gamma_{i}=1$, $\forall i=1...n$, $\hat{\mathbf{m}_{i}}=[\mathbf{m}_{i}^{T},\mu_{i}]^{T}$, and $\hat{\mathbf{n}_{i}}=[\mathbf{n}_{i}^{T},\nu_{i}]^{T}$ the estimation problem of (3) and (1) can be represented by equations (5) and (6):
@@
In case of $\alpha_{i}=1$, $\beta_{i}=1$, $\gamma_{i}=1$, $\mu_{i}=0$, and $\nu_{i}=0$, we have tranformed the affine transformation into a linear form.
Figure 5: Simulation result, mean errors $\bar{e}_{x}$ and $\bar{e}_{y}$ for the augmented implementation of Gleser and Watson compared to the augmented Alg. 1 as a function of the standard deviation $\sigma$.
@@
III-A Solution proposed by Gleser and Watson
@@
Let $\boldsymbol{\lambda}$ be the diagonal matrix of the $p$ largest eigenvalues and $\mathbf{U}$ be the matrix of all corresponding eigenvectors of $\mathbf{X}^{T}\mathbf{X}+\mathbf{Y}^{T}\mathbf{Y}$, then the estimates of $\boldsymbol{\Theta}$ and $\mathbf{B}$ are given by (9) and (10):
@@ III-B Multivariate Regression Approach
Important for an implementation on low-power IoT-devices is a balancing of the computation complexity of the algorithms. Therefore, we also investigate an alternative solution in this section. The solution of Gleser and Watson in Lemma 1 needs an eigenvector decomposition. We are interested in an estimation of the AT from $\mathbf{X}$ to $\mathbf{Y}$. Let $\mathbf{r}_{i}=\hat{\mathbf{n}}_{i}-\mathbf{B}\hat{\mathbf{m}}_{i}$, we can rewrite (6) to (11):
@@
This leads to the idea of of a simple least squares estimate [1]. In general, the sensor measurement is noisy. However, in cases of high quality sensors, the noise power is very low. Hence, a simple regression approach has advantages, e.g., lower computational complexity and thus energy consumption.
As observed in [1], if $\boldsymbol{\Theta}\boldsymbol{\Theta}^{T}$ is non-singluar, the minimum over $\mathbf{B}$ of $\mathrm{tr}[(\mathbf{Y}-\mathbf{B}\boldsymbol{\Theta})(\mathbf{Y}-\mathbf{B}\boldsymbol{\Theta})^{T}]$ can be obtained based on the results of multivariate regression $\mathbf{B}=\mathbf{Y}\boldsymbol{\Theta}^{T}[\boldsymbol{\Theta}\boldsymbol{\Theta}^{T}]^{-1}$. This solution corresponds to the search for $\mathbf{B}$ with $\mathbf{B}\boldsymbol{\Theta}=\mathbf{Y}$. The resulting matrix $\mathbf{B}$ is obtained by multiplying $\mathbf{Y}$ by the pseudoinverse of $\boldsymbol{\Theta}$ (i. e., the Moore-Penrose inverse), i.e. $\mathbf{B}=\mathbf{Y}\boldsymbol{\Theta}^{T}[\boldsymbol{\Theta}\boldsymbol{\Theta}^{T}]^{-1}$. This solution is the least squares solution. Interestingly, the stationary points of (8) are $\boldsymbol{\Theta}=\mathbf{X}$ and $\mathbf{Y}=\mathbf{B}\boldsymbol{\Theta}$.
@@
Algorithm 1 Augmented implementation of Gleser and Watson (denoised)
@@
1:  $[\mathbf{U},\boldsymbol{\Lambda}]=\mathrm{eigs}(\mathbf{X}^{T}\mathbf{X}+\mathbf{Y}^{T}\mathbf{Y},p)$
2:  $\boldsymbol{\Theta}^{T}=\mathbf{U}\mathbf{U}^{T}\mathbf{X}^{T}$
3:  $\boldsymbol{\Theta}(p,:)=\mathbf{1}^{T}$
4:  $\mathbf{B}=\mathbf{Y}\boldsymbol{\Theta}^{T}(\boldsymbol{\Theta}\boldsymbol{\Theta}^{T})^{-1}$
5:  $\tilde{\mathbf{A}}=\mathbf{B}(1:p-1,1:p-1)$
6:  $\tilde{\mathbf{b}}=\mathbf{b}(1:p-1,p)$
7:  $\boldsymbol{\Theta}_{E}=\boldsymbol{\Theta}(1:p-1,:)$
Algorithm 1 is the augmented implementation of Gleser and Watson. In line $2$ we calculate the $p$ largest eigenvalues $\boldsymbol{\Lambda}$ according to Lemma 1, and use the associated eigenvectors to estimate $\boldsymbol{\Theta}$, the origins of $\mathbf{X}$, in line $3$. The variables $\gamma_{i}$ should satisfy $\gamma_{i}=1$, $\forall i=1,...,n$. Therefore, after step $2$ of Algorithm 1 set $\gamma_{i}=1$, $\forall i=1,...,n$. The resulting $\boldsymbol{\Theta}$ augmentation line contains many values close to $1$, so we set those to the value $1$ in line $4$ to suppress the small eigenvalues impact on $\boldsymbol{\Theta}$ with low noise power [25]. Lines $5$ to $7$ calculate and extract the estimated AT parameters and line $8$ removes the augmentation to format the origin estimation.
Algorithm 2 corresponds to the derivation of the simple least squares according to Proposition 1. Since our focus lies not in the estimation of the origins $\boldsymbol{\Theta}$, we can directly calculate and extract the estimatation parameters in lines $2$ to $5$. Although we cannot achieve improved complexity, it should be possible to achieve lower energy consumption through an improved number of operations, since Alg. 2 skips step 1 and step 2 of Alg. 1. Furthermore, the accuracy is very close to the results of Alg. 3 (cf. Table I and Table II), which, however, requires a similar number of operations as Alg. 1.
@@
1:  $\mathbf{B}=\mathbf{Y}\mathbf{X}^{T}(\mathbf{X}\mathbf{X}^{T})^{-1}$
2:  $\tilde{\mathbf{A}}=\mathbf{B}(1:p-1,1:p-1)$
3:  $\tilde{\mathbf{b}}=\mathbf{b}(1:p-1,p)$
4:  $\boldsymbol{\Theta}_{E}=\mathbf{X}(1:p-1,:)$
The simple least squares solution of Alg. 2 and the solution of Gleser and Watson in Alg. 1 can be combined to the improved Alg. 3 by using the following observations.
III-C Hybrid Approach